OpenStudy (anonymous):

the price of 9-volt batteries is increasing according to the function below. where t is year after anuary 1,1980. during what year will the price reach 3 dollars? p(t)=1.1*e^0.047t *=multiply

OpenStudy (btaylor):

\[3=1.1e^{0.047t}\]\[2.7272=e^{0.047t}\]\[\ln(2.7272)=0.047t \ln(e)\]\[\ln(2.7272)=0.047t\]\[t=\frac{\ln(2.7272)}{0.047} = 21.346 years\]. So 21 years after 1980 would be 2001

OpenStudy (anonymous):

@BTaylor HOLY CRAP YOURE A GENIUS THAN YOU SOOOO MUCHH sorry i didnt realize caps was on...